OpenStudy (anonymous):

An amount of $34,000 is borrowed for 13 years at 6.25% interest, compounded annually. If the loan is paid in full at the end of that period, how much must be paid back?

OpenStudy (anonymous):

@jackthegreatest

OpenStudy (solomonzelman):

the formula for compound interest is: \(\LARGE\color{blue}{ A=P\left(\begin{matrix} 1+\frac{\LARGE r}{\LARGE n} \\ \end{matrix}\right)^{n~\times~t} }\) where: \(\LARGE\color{blue}{ A}\) is the value you get in the future: \(\LARGE\color{blue}{ P }\) is the initial investment (or amount) \(\LARGE\color{blue}{ r }\) is a percent rate (in decimal). \(\LARGE\color{blue}{ n }\) is the number of times the interest is compound each year. \(\LARGE\color{blue}{ t }\) is the number of years

OpenStudy (solomonzelman):

in your case: \(\LARGE\color{green}{ A}\) is what you want to find \(\LARGE\color{blue}{ P }\) is 34,000 \(\LARGE\color{blue}{ r }\) is 0.0625 (converted the % by dividing by 100) \(\LARGE\color{blue}{ n }\) is 1. (because it is compounded annually, as the problem tells you) \(\LARGE\color{blue}{ t }\) is 13 (it is given to be so).

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ A=34,000\left(\begin{matrix} 1+\frac{\LARGE 0.0625}{\LARGE 1} \\ \end{matrix}\right)^{1~\times~13} }\) I plugged the numbers for you and leaving it for you to calculate. good luck!